# Leakage and Interpretability in Concept-Based Models

### v1

@@ 5 Concepts-task and interconcept leakage scores
Leakage is information-theoretic in nature and as such, it can be captured appropriately only by metrics based on quantities from information theory. As discussed in Section 3, the leaked information and the corresponding additional structure in the learnt concept distributions can be measured as modifications in the ground-truth interconcept and concepts-task MIs respectively. This motivates the definition of the following set of scores based on information theory to quantify leakage and the interpretability of concept-based models, and overcome the limitations of the existing measures.
@@
$CTL_{i}(\hat{c},c,y)=\left|\frac{I(\hat{c}_{i},y)}{H(y)}-\frac{I(c_{i},y)}{H(y)}\right|\,.$ (6)
This score quantifies the additional information that the learnt concept $i$ encodes about the task label with respect to ground-truth. MIs are appropriately normalised by the entropy of the class label to ensure $0\leq CTL_{i}\leq 1$. We further define the average concepts-task leakage score
@@
as the average extra information about the task stored into the learnt concepts.
@@
$ICL_{ij}(\hat{c},c)=\left|\frac{I(\hat{c}_{i},\hat{c}_{j})}{\sqrt{H(\hat{c}_{i})\,H(\hat{c}_{j})}}-\frac{I(c_{i},c_{j})}{\sqrt{H(c_{i})\,H(c_{j})}}\right|\,,$ (8)
which measures the additional predictivity of the learnt concept $i$ for concept $j$ with respect to ground-truth. The normalisation by the entropies also in this case ensures $0\leq ICL_{ij}\leq 1$, and note that $ICL_{ii}=0$ since $I(z,z)=H(z)$. We also define the per-concept and average interconcept leakage scores
@@
The $CTL$ and $ICL$ scores are meant to summarise complementary aspects of the leakage present in a given concept-based model, as well as its overall interpretability. We define the following criterion to detect leakage using the proposed scores:
@@
The $CTL_{i}$, $ICL_{ij}$ and $ICL_{i}$ measures convey more detailed information about each type of leakage at the level of single concepts. This is particularly useful for model design and when performing interventions (as they essentially indicate the risk of intervening on each concept). The framework above can be applied to concept-based models with a range of concept encodings, including binaries, soft probabilities, logits and vector representations. Alternative normalisations of MI can be used, although we found those adopted in (6) and (8) to be particularly transparent from an information theory perspective, and robust in the experiments carried out. We estimate MI and entropy using the KSG estimator (Kraskov et al., 2004), based on $k$-nearest neighbour statistics.
@@ 6 Robustness of leakage scores
$c_{acc}$ $c_{\text{F1}}$ $c_{\text{AUC}}$ $y_{acc}$ $y_{\text{F1}}$ $y_{\text{AUC}}$ $S_{(int)}(\downarrow)$
@@
Table 1: Concept and task evaluation scores, as well as intervention scores of pairs of soft and logit CBMs with different levels of leakage (see Figure 4).
@@
For our experiments we consider three synthetic datasets, TabularToy($\delta$), dSprites($\gamma$) and 3dshapes($\gamma$), presented in Espinosa Zarlenga et al. (2023a). TabularToy($\delta$) is a binary-class tabular dataset based on the dataset from Mahinpei et al. (2021), while dSprites($\gamma$) and 3dshapes($\gamma$) are multi-class image-based datasets building upon dSprites (Matthey et al., 2017) and 3dshapes (Burgess and Kim, 2018) respectively. The use of synthetic datasets enables us to (i) modify the functional dependence $y=f(c)$ and hence tune the ground-truth concepts-task MI, and furthermore (ii) tune the ground-truth interconcept MI via the parameters $\delta\in(0,1)$ and $\gamma$ ($\in\{0,\dots,4\}$ in dSprites and $\in\{0,\dots,5\}$ in 3dshapes). See Appendix A for more details on the experimental setup.
@@
We consider pairs of models encoding different levels of leakage as assessed by intervention performance (Figure 4). The models in each pair were chosen to have essentially identical concepts and task evaluation scores (Table 1), evidencing the severe limitations of these metrics in capturing leakage. In all cases the CTL and ICL scores correctly detect the different amounts of leakage according to the Leakage Criterion, with no level of uncertainty. On the contrary, we note as a general behaviour that the large confidence intervals of the OIS prevent drawing any statistically significant conclusion on leakage. Appendix C presents the concept-wise leakage scores for these models and illustrates how they provide more fine-grained information at the level of individual concepts.
@@
Figure 5 shows the leakage scores for hard models trained on a number of datasets in both low and high regimes of ground-truth interconcept MI. By construction hard CBMs have vanishing leakage (Koh et al., 2020; Kazhdan et al., 2021; Margeloiu et al., 2021), and the CTL and ICL scores are in all cases compatible with zero, regardless of the interconcept MI. The OIS is instead non-vanishing in all experiments, failing to meet a fundamental requirement of leakage metrics. We further note that for each dataset its estimated values on hard CBMs are comparable to those on the models shown in Figure 4, which instead encode a significant amount of leakage. This represents further evidence of the limitations of the OIS in distinguishing interpretable from uninterpretable concept-based models.
Figure 5: Leakage scores evaluated on hard CBMs on datasets with high and low ground-truth interconcept MI. The 95% confidence intervals are obtained from a 5-fold training on each dataset.
@@
Pearson $r$ $-\log_{10}(p)$ Pearson $r$ $-\log_{10}(p)$ Pearson $r$ $-\log_{10}(p)$
TT($0.25$) 0.75 $\pm$ 0.05 2.6 $\pm$ 0.4 0.69 $\pm$ 0.07 2.1 $\pm$ 0.5 0.49 $\pm$ 0.11 1.1 $\pm$ 0.4
dS(00) 0.83 $\pm$ 0.06 3.6 $\pm$ 0.8 0.68 $\pm$ 0.07 2.0 $\pm$ 0.5 0.69 $\pm$ 0.07 2.2 $\pm$ 0.5
3ds(00) 0.84 $\pm$ 0.02 4.0 $\pm$ 0.3 0.86 $\pm$ 0.03 4.4 $\pm$ 0.5 0.81 $\pm$ 0.04 3.9 $\pm$ 0.7
TT($0.75$) 0.55 $\pm$ 0.05 1.3 $\pm$ 0.2 0.55 $\pm$ 0.12 1.4 $\pm$ 0.5 0.27 $\pm$ 0.13 0.5 $\pm$ 0.3
dS($4$) 0.70 $\pm$ 0.09 2.2 $\pm$ 0.6 0.45 $\pm$ 0.08 1.0 $\pm$ 0.3 0.55 $\pm$ 0.10 1.3 $\pm$ 0.5
3ds($5$) 0.75 $\pm$ 0.04 2.9 $\pm$ 0.4 0.24 $\pm$ 0.15 0.5 $\pm$ 0.5 0.70 $\pm$ 0.07 2.5 $\pm$ 0.6
Table 2: Pearson $r$ coefficients and corresponding $p$-values measuring the correlations of CTL, ICL and OIS metrics against $S_{(int)}$. To better capture the scores distributions resulting from MC sampling, the displayed uncertainties are standard deviations. Significant and highly significant correlations correspond to $-\log_{10}(p)\simeq 1.3$ and $2.0$ respectively.
@@
We train a number of soft and logit CBMs on each dataset in both high and low regimes of ground-truth interconcept MI and with different levels of concept supervision (see Appendix D for details on this experiment). The evaluation of the OIS, ICL and CTL scores is repeated 5 times for each individual model yielding an uncertainty on their value. We then Monte-Carlo (MC) sample 10K times from the uncertainty distributions of each score (assuming normality), and for each sample we compute the Pearson $r$ coefficient between each score and the intervention score $S_{(int)}$. The resulting distribution of Pearson $r$ values between each leakage score and intervention performance across models and for each dataset provides an estimate of its uncertainty, while the corresponding distribution of $p$-values provide information about the significance of correlation.
Table 2 reports the mean and standard deviation of the resulting Pearson $r$ and $p$-value distributions. The proposed CTL and ICL scores highly correlate with intervention performance and they systematically outperform OIS across datasets and robustly to ground-truth interconcept MI. This analysis also quantifies the importance of concepts-task and interconcept leakage in each case. CTL generally appears as the prominent form of leakage over ICL. Its correlation with intervention performance is typically stronger than ICL, and is moreover significant in all datasets and highly significant in 5 out of 6 datasets. This behaviour reflects the optimization pressure favouring the storage of additional task-relevant information in the learnt concepts rather than information about the other concepts, to achieve a better task performance. Nevertheless, these results are evidence that also interconcept leakage systematically arises during model training, in particular displaying correlations of comparable magnitude to concept-task leakage in 3 out of 6 datasets.
@@ 7 Causes of leakage
The proposed information-theoretic leakage scores are robust and sensitive and can thus be used to identify the most common causes of leakage and assess their impact on interpretability.
A first effect that directly impacts the interpretability of a concept-based model is the concept encoder being misspecified with respect to the target concept distribution. When such a misspecification is significant, the model is essentially not capable of learning concepts and their relations. This may favour the emergence of additional issues, such as non-locality (Raman et al., 2023; Huang et al., 2024). Concept-based architectures enforcing accurate learning of interconcept relations have been recently proposed (Kim et al., 2023; Vandenhirtz et al., 2024; Xu et al., 2024; Kim et al., 2025). We recognise however that this problem is significantly use-case specific and it is usually sufficient to opt for a concept encoder with an improved architecture to solve it. In the following experiments we will thus consider concept encoders that are sufficiently well-specified.
@@
When $\lambda$ is set to be small, during training the model is encouraged to achieve better task performance at the cost of poor concept learning and high leakage. This issue was originally observed in Koh et al. (2020), where poor intervention performance was found when training models with low $\lambda$. To precisely quantify this effect, we train soft and logit CBMs with low, intermediate and high values of $\lambda$ on datasets with both low and high interconcept MI. The resulting leakage scores displayed in Figure 6 confirm that leakage decreases as we enhance concept supervision.
However, as detailed in Appendix E, leakage does not fall to zero as one keeps increasing $\lambda$ – the leakage scores reach a non-vanishing minimal value, specific to each model class. At higher values $\lambda\gtrsim 10$, models focus excessively on concept learning, and the task objective may be poorly learnt. This results in a drop in task performance or an increase in leakage. Each model class has thus an associated optimal range of $\lambda\in(\lambda_{min},\lambda_{max})$ where the minimal amount of leakage is attained, below which models exhibit proportionally higher leakage and above which training often fails. As evidenced by our results, such an optimal interval can be identified using the CTL and ICL scores.
@@
When the chosen concept representation is significantly more expressive than the annotated concept representation, during training the model is generally encouraged to misuse the surplus expressivity to store additional information into concepts to ease task predictions. This is evidenced in Figure 6 where annotated concepts are binary, while learnt concepts are soft probabilities and logits, and as such over-expressive for the setup. We note that raising $\lambda$ is less effective at decreasing leakage when concepts representations are more expressive, as in the logit models, and furthermore the minimal attainable leakage is significantly higher for such models over soft CBMs. This behaviour casts doubts on the advantages of vector concept encodings when ground-truth concepts are binary, and we will see in Section 8 how these concerns impact and are substantiated in CEMs. Finally, we observe that models with low $\lambda$ generally have a comparable level of high leakage regardless of the concept encoding.
@@
When the set of annotated concepts is incomplete for the task, meaning that it is not sufficiently predictive, models tend to encode the information about the missing concepts (or generally additional information that may be required) into the learnt concepts to nonetheless achieve a high task performance. This translates into sizeable leakage and a loss of interpretability. This issue has been discussed in previous works (Kazhdan et al., 2021; Margeloiu et al., 2021; Mahinpei et al., 2021; Havasi et al., 2022).
To accurately assess this phenomenon using our information-theoretic scores, we train soft CBMs with low, intermediate and high levels of concept supervision on a range of datasets with roughly one third of concepts removed (see Appendix A for more details on which concepts are removed for each dataset). This results in a sizeable decrease of the reference task accuracy $y_{acc}^{(k)}$ of the final head trained on the ground-truth concepts introduced in equation (5) (Table 3, left). The leakage scores of the trained models are displayed in Figure 7, along with those of models with complete sets of concepts for comparison. Their concepts and task accuracies are presented in Appendix F.
According to the Leakage Criterion, we detect more leakage in 9/9 model classes trained on incomplete sets of concepts. At low and intermediate $\lambda$, there is considerably more leakage in models trained on an incomplete set. At high lambda, CTL scores are typically comparable in models with incomplete and complete concept sets, while ICL remains significantly higher in case of an incomplete concept set.
@@
Table 3: Former two columns: baseline $y_{acc}^{(k)}$ task accuracies of final heads trained on datasets with complete and incomplete sets of concepts. Latter two columns: baseline $y_{acc}^{(k)}$ task accuracies of a linear head on the linear (well-specified) and non-linear (misspecified) ground-truth task labels. We report 95% confidence intervals over 5-fold training.
@@
If the final head is not flexible enough to learn sufficiently well the ground-truth $y=f(c)$ dependence, during training the model tends to store the necessary dependences into the learnt concepts, which get deformed to encode additional information about the task and the other concepts. In real-world applications a common choice for the final head of a CBM is a linear layer, which being itself interpretable, allows in principle for a full model explainability. In reality the task is however very often described by a highly non-linear function of the concepts, and a linear layer can be severely misspecified in such cases.
As an example showcasing the general effects of final head misspecification as well as representing a potential shortcoming of a considerable fraction of existing CBMs, we consider soft CBMs with a linear head trained on datasets where the ground-truth task is either a linear or non-linear function of the concepts. To achieve that, we modify the functional dependences in TabularToy(0.25), dSprites(0) and 3dshapes(0) by adding non-linear terms to the original tasks – see Appendix A.1. As a baseline measuring the final head misspecification, for each dataset we train a separate linear head on the ground-truth concepts. Table 3 shows the resulting decrease in task performance in terms of the reference task accuracy $y^{(k)}_{acc}$. The corresponding leakage scores are displayed in Figure 8 and their concepts and task accuracies are presented in Appendix F. According to the Leakage Criterion, a misspecified final head causes a higher leakage in 7/9 model classes. Note in particular that high concept supervision is generally not sufficient to compensate for the additional leakage induced by the final head misspecification.

### v2

@@ 5 Concepts-task and interconcept leakage scores
Information leakage is information-theoretic in nature and, as such, it is most appropriately captured by metrics based on quantities from information theory. As discussed in Section 3, the leaked information and the corresponding additional structure in the learnt concept distributions can be measured as modifications in the ground-truth interconcept and concepts-task MIs respectively. This motivates the definition of the following set of scores based on information theory to quantify leakage and the interpretability of concept-based models:
@@
$CTL_{i}(\hat{c},c,y)=\max\left(0,\frac{I(\hat{c}_{i},y)}{H(y)}-\frac{I(c_{i},y)}{H(y)}\right).$ (6)
This score quantifies the additional information that the learnt concept $i$ encodes about the task label with respect to the ground-truth. MIs are appropriately normalised by the entropy of the class label to ensure $0\leq CTL_{i}\leq 1$. When the difference inside the brackets is negative, the learnt concepts are less predictive of the task than expected from the ground truth, indicating poor concept learning rather than leakage. We further define the average concepts-task leakage score as
@@
as the average extra information about the task stored into the learnt concepts. Note that a weighted average over concepts can be adopted in cases where concept importances are non-uniform.
@@
$ICL_{ij}(\hat{c},c)=\max\left(0,\frac{I(\hat{c}_{i},\hat{c}_{j})}{\sqrt{H(\hat{c}_{i})\,H(\hat{c}_{j})}}-\frac{I(c_{i},c_{j})}{\sqrt{H(c_{i})\,H(c_{j})}}\right),$ (8)
which measures the additional predictivity of the learnt concept $i$ for concept $j$ with respect to ground-truth. Again, normalisation by the entropies ensures that $0\leq ICL_{ij}\leq 1$. Note that $ICL_{ii}=0$ since $I(z,z)=H(z)$. We also define the per-concept and average interconcept leakage scores as
@@
The $CTL$ and $ICL$ scores are meant to summarise complementary aspects of the leakage present in a given concept-based model, as well as its overall interpretability. Based on these measures, we define the following criterion to detect leakage:
@@
These measures also convey more detailed information about each type of leakage at the level of single concepts. This is particularly useful for model design and when performing interventions (as they essentially indicate the risk of intervening on each concept). The framework above can be applied to concept-based models with a range of concept encodings, including binaries, soft probabilities, logits and vector representations. Alternative normalisations of MI can be used, although we found those adopted in (6) and (8) to be particularly transparent from an information theory perspective, and robust in the experiments we carried out. We estimate MI and entropy using the Kraskov-Stögbauer-Grassberger (KSG) estimator (Kraskov et al., 2004), based on $k$-nearest neighbour statistics.
@@ 6 Robustness of leakage scores
$c_{acc}$ $c_{\text{F1}}$ $c_{\text{AUC}}$ $y_{acc}$ $y_{\text{F1}}$ $y_{\text{AUC}}$ $S_{int}(\downarrow)$
@@
Table 1: Concept and task evaluation scores, and intervention scores for pairs of soft and logit CBMs with different levels of leakage (see Figure 4).
@@
For our experiments we considered three synthetic datasets, TabularToy($\delta$), dSprites($\gamma$) and 3dshapes($\gamma$), presented in Espinosa Zarlenga et al. (2023a). TabularToy($\delta$) is a binary-class tabular dataset based on the dataset from Mahinpei et al. (2021), while dSprites($\gamma$) and 3dshapes($\gamma$) are multi-class image-based datasets building upon dSprites (Matthey et al., 2017) and 3dshapes (Burgess and Kim, 2018) respectively. The use of synthetic datasets enables us to (i) modify the functional dependence $y=f(c)$ and hence tune the ground-truth concepts-task MI, and (ii) tune the ground-truth interconcept MI via the parameters $\delta\in(0,1)$ and $\gamma$ ($\in\{0,\dots,4\}$ in dSprites and $\in\{0,\dots,5\}$ in 3dshapes), an essential aspect to assess interpretability (Heidemann et al. (2023)). See Appendix A for more details of these experiments.
@@
We consider pairs of models encoding different levels of leakage as assessed by intervention performance (Figure 4). The models in each pair were chosen to have essentially identical concepts and task evaluation scores (Table 1), evidencing the limitations of these metrics in capturing leakage. In all cases the CTL and ICL scores correctly detect the different amounts of leakage according to the Leakage Criterion, with minimal uncertainty. In contrast, the OIS has large confidence intervals and so does not allow statistically significant conclusion on leakage to be drawn. Appendix C presents the concept-wise leakage scores for these models and illustrates how they provide more fine-grained information at the level of individual concepts.
@@
Figure 5 shows the leakage scores for hard models trained on a number of datasets in both low and high regimes of ground-truth interconcept MI. By construction hard CBMs have vanishing leakage (Koh et al., 2020; Kazhdan et al., 2021; Margeloiu et al., 2021). In all cases, the CTL and ICL scores do not deviate significantly from zero, regardless of the interconcept MI. In contrast, the OIS is non-vanishing in all cases, indicating that it does not meet a fundamental requirement as a measure of leakage. We further note that for each dataset the estimated OIS for hard CBMs are comparable to those of the models shown in Figure 4, which exhibit a significant amount of leakage. Thus, in these cases OIS does not distinguish interpretable from uninterpretable concept-based models.
Figure 5: Leakage scores evaluated on hard CBMs on datasets with high and low ground-truth interconcept MI. The 95% confidence intervals are obtained from 5-fold training on each dataset.
@@
Pearson $r$ $p$-value Pearson $r$ $p$-value Pearson $r$ $p$-value
TT($0.25$) 0.75 $\bm{3.4\times 10^{-3}}$ 0.69 $1.4\times 10^{-2}$ 0.49 $1.2\times 10^{-1}$
dS(00) 0.83 $\bm{1.1\times 10^{-3}}$ 0.68 $1.5\times 10^{-2}$ 0.70 $1.1\times 10^{-2}$
3ds(00) 0.84 $4.2\times 10^{-5}$ 0.86 $\bm{2.4\times 10^{-5}}$ 0.82    $\!\,\,\,2.0\times 10^{-4}$
TT($0.75$) 0.54 $\bm{5.9\times 10^{-2}}$ 0.51 $1.1\times 10^{-1}$ 0.27 $4.1\times 10^{-1}$
dS($4$) 0.67 $\bm{1.9\times 10^{-2}}$ 0.48 $1.2\times 10^{-1}$ 0.55 $7.4\times 10^{-2}$
3ds($5$) 0.79 $\bm{3.9\times 10^{-4}}$ 0.69 $3.5\times 10^{-2}$ 0.70 $6.5\times 10^{-3}$
Table 2: Pearson $r$ coefficients and corresponding $p$-values measuring the correlations of CTL, ICL and OIS metrics against $S_{int}$.
@@
We trained a number of soft and logit CBMs on each dataset in both high and low regimes of ground-truth interconcept MI and with different levels of concept supervision (see Appendix D for details on this experiment). The evaluation of the OIS, ICL and CTL scores was repeated 5 times for each individual model to provide estimates of uncertainty. We then took 10,000 Monte-Carlo (MC) samples from the resulting distributions of each score (assuming normality), and for each sample we computed the Pearson $r$ coefficient between each score and the intervention score $S_{int}$. Pooling using Rubin’s rule (see Rubin, 2004; Barnard and Rubin, 1999), we obtain a single best estimate for the Pearson $r$ and its associated $p$-value for each leakage score and dataset, which are reported in Table 2.
The proposed CTL and ICL scores strongly correlate with intervention performance, they systematically outperform OIS across all datasets and correspond robustly to ground-truth interconcept MI. This analysis also quantifies the importance of concepts-task and interconcept leakage in each case. CTL generally appears as the prominent form of leakage over ICL. Its correlation with intervention performance is typically stronger than ICL, and is, moreover, significant in all datasets and highly significant in 5 out of 6 datasets. This behaviour suggests that optimization favours the storage of additional task-relevant information in the learnt concepts rather than information about the other concepts. Nevertheless, these results also provide evidence that interconcept leakage systematically arises during model training; indeed, correlations of comparable magnitude to concept-task leakage were observed in the experiments on 3 out of 6 datasets.
@@ 7 Causes of leakage
The previous results indicate that our proposed information-theoretic leakage scores are robust and sensitive and can therefore be used to identify the most common causes of leakage and assess their impact on interpretability.
A first effect that directly impacts the interpretability of a concept-based model is the concept encoder being misspecified with respect to the target concept distribution. When such a misspecification is significant, the model is essentially not capable of learning concepts and their relations. This may result in additional issues, such as non-locality (Raman et al., 2023; Huang et al., 2024). To address this issue, concept-based architectures that enforce accurate learning of interconcept relations have been recently proposed (Kim et al., 2023; Vandenhirtz et al., 2024; Xu et al., 2024; Kim et al., 2025). We recognise however that this problem is strongly use-case specific and it is usually sufficient to opt for a concept encoder with an improved architecture to solve it. In the following experiments we will thus consider concept encoders that are sufficiently well-specified, as relevant for real-world high-risk scenarios where the aim is both high concept performance and the absence of leakage.
@@
When $\lambda$ is set to be small in the loss (3) during training, the model will achieve better task performance at the cost of poor concept learning and high leakage. This issue was originally observed in Koh et al. (2020), where poor intervention performance was found when training models with low $\lambda$. To precisely quantify this effect, we trained soft and logit CBMs with low, intermediate and high values of $\lambda$ on datasets with both low and high interconcept MI. The resulting leakage scores, displayed in Figure 6, confirm that leakage decreases as concept supervision is increased.
However, as detailed in Appendix E, leakage does not fall to zero as $\lambda$ is increased – instead, the leakage scores reach a non-vanishing minimal value, specific to each model class. At higher values (here, $\lambda\gtrsim 10$), models focus excessively on concept learning, and the task objective may be poorly learnt. This results in a drop in task performance or an increase in leakage. Each model class thus has an associated optimal range of $\lambda\in(\lambda_{min},\lambda_{max})$ where the minimal amount of leakage is attained, below which models exhibit proportionally higher leakage and above which training often fails. As evidenced by our results, such an optimal interval can be identified using the CTL and ICL scores.
@@
When the chosen concept representation is significantly more expressive than the annotated concept representation, the model will generally misuse the surplus expressivity to store additional information in the concepts to aid task prediction. This phenomenon is illustrated in Figure 6 where annotated concepts are binary, while learnt concepts are soft probabilities and logits, and as such over-expressive for the setup. We note that raising $\lambda$ is less effective at decreasing leakage when concepts representations are more expressive, as in the logit models, and furthermore the minimal attainable leakage is significantly higher for such models over soft CBMs. This behaviour casts doubt on the advantages of vector concept encodings when ground-truth concepts are binary. As we will show in Section 8 these concerns are substantiated in CEMs. Finally, we observe that models with low $\lambda$ generally have a comparable level of high leakage regardless of the concept encoding.
@@
When the set of annotated concepts is incomplete for the task – meaning that it is not sufficiently predictive – models tend to encode information about the missing concepts (or, more generally, any additional information that may be required) into the learnt concepts to nonetheless achieve a high task performance. This typically translates into sizeable leakage and a loss of interpretability. This issue has been discussed in previous work (Kazhdan et al., 2021; Margeloiu et al., 2021; Mahinpei et al., 2021; Havasi et al., 2022).
To accurately assess this phenomenon using our information-theoretic scores, we trained soft CBMs with low, intermediate and high levels of concept supervision on a range of datasets with roughly one third of concepts removed (see Appendix A.1 for more details on which concepts are removed for each dataset). This resulted in a sizeable decrease of the reference task accuracy $y_{acc}^{(k)}$ of the final head trained on the ground-truth concepts introduced in equation (5) (Table 3, left). The leakage scores of the trained models are displayed in Figure 7, along with those of models with complete sets of concepts for comparison. The relevant concepts and task accuracies are presented in Appendix F.
According to the Leakage Criterion and as expected, we detected more leakage in 8/9 model classes trained on incomplete sets of concepts. At low and intermediate values of $\lambda$, there is significantly more leakage in models trained on an incomplete set. At high lambda, CTL scores are typically comparable in models with incomplete and complete concept sets, while ICL remains significantly higher in case of an incomplete concept set.
@@
Table 3: First two columns: baseline $y_{acc}^{(k)}$ task accuracies of final heads trained on datasets with complete and incomplete sets of concepts. Second two columns: baseline $y_{acc}^{(k)}$ task accuracies of a linear head on the linear (well-specified) and non-linear (misspecified) ground-truth task labels. We report 95% confidence intervals over 5-fold training.
@@
If the final head is not flexible enough to learn the ground-truth $y=f(c)$ dependence, during training then the model tends to store the necessary dependences into the learnt concepts, which are thereby deformed to encode additional information about the task and the other concepts. In many practical applications a common choice for the final head of a CBM is a linear layer which in principle allows for full model explainability. In reality, many tasks are best described by a non-linear function of the concepts, and a linear layer can be misspecified in such cases.
To explore the general effects of final head misspecification we considered soft CBMs with a linear head trained on datasets where the ground-truth task is either a linear or non-linear function of the concepts. To do so we modified the functional dependences in TabularToy(0.25), dSprites(0) and 3dshapes(0) by adding non-linear terms to the original tasks (see Appendix A.1 for details). As a baseline measuring the final head misspecification, for each dataset we trained a separate linear head on the ground-truth concepts. Table 3 shows the resulting decrease in task performance in terms of the reference task accuracy $y^{(k)}_{acc}$. The corresponding leakage scores are displayed in Figure 8 and their concepts and task accuracies are presented in Appendix F. According to the Leakage Criterion, a misspecified final head causes a higher leakage in 7/9 model classes. Moreover, high concept supervision is generally not sufficient to compensate for the additional leakage induced by the final head misspecification.